# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v2

@@ Appendix C Estimators of Thresholds
For every constraint $i\in[m]$, state $s$, action $a$, step $h\in[H]$ and episode $t\in[T]$, let $(s_{h}^{l},a_{h}^{l})$ denote the state-action pair visited in episode $l$ at step $h$; $(s_{h}^{l},a_{h}^{l},s_{h+1}^{l})$ represents the state-action pair $(s_{h}^{l},a_{h}^{l})$ is visited and the environment evolves to next state $s_{h+1}^{l}$ at step $h$ in episode $l$; $\mathbf{1}_{X}$ is the indicator function of $X$; and $N_{h}^{t}(s,a)=\sum_{l=1}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ is the total number of visits to the pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ at step $h$ before episode $t\in[T]$. First, we give the empirical averages of the thresholds as follows:
[…]
@@
$\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{t}(s,a),$
where $\iota_{h}^{t}(s,a):=\sqrt{\frac{H^{2}\ln\left(\frac{2}{\delta}\right)}{2N_{h}^{t}(s,a)}}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\geq\frac{c}{N_{h}^{t}(s,a)}\right]\leq 2\exp\left(-\frac{2c^{2}}{H^{2}N_{h}^{t}(s,a)}\right).$
Setting $\delta=2\exp\left(-\frac{2c^{2}}{H^{2}N_{h}^{t}(s,a)}\right)$ and solving to find a proper value of $c$ gives the result for the threshold function. ∎
@@
$\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\zeta_{h}^{t}(s,a),$
where $\zeta_{h}^{t}(s,a):=\min\left\{H,\sqrt{\frac{4H^{2}\ln(mSAHT/\delta)}{\max\{1,N_{h}^{t}(s,a)\}}}\right\}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{t}(s,a)\right]\geq 1-\delta^{\prime}.$
@@
$\mathbb{P}\left[\bigcap_{s,a,m,h,t}\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{t}(s,a)\right\}\right].$
@@
$\mathbb{P}\left[\bigcap_{s,a,m,h,t}\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{t}(s,a)\right\}\right]$
$=$ $1-\mathbb{P}\left[\bigcup_{s,a,m,h,t}\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{t}(s,a)\right\}^{c}\right]$
$=$ $1-\sum_{s,a,m,h,t}\mathbb{P}\left[\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{t}(s,a)\right\}^{c}\right]$
@@
where the inequality holds by union bound. Noticing that $\alpha_{i,h}^{t}(s,a)\leq H$, we substitute $\delta^{\prime}$ with $\delta^{\prime}=\frac{\delta}{mSAHT}$ in $\iota_{t}(s,a)$ with an additional union bound over the possible values of $N_{h}^{t}(s,a)$. This yields $\zeta_{t}(s,a)$, and concludes the proof. ∎
@@
$\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{(W_{t}),t}(s,a),$
where $\iota_{h}^{(W_{t}),t}(s,a)=\sqrt{\frac{H^{2}\ln\left(\frac{2}{\delta}\right)}{2\max\{1,N_{h}^{(W_{t}),t}(s,a)\}}}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}(s,a)\right|\geq\epsilon\right]\leq 2\exp\left(-\frac{2\epsilon^{2}N_{h}^{(W_{t}),t}(s,a)}{H^{2}}\right).$
Setting $\delta=2\exp\left(-\frac{2\epsilon^{2}N_{h}^{(W_{t}),t}(s,a)}{H^{2}}\right)$ and solving for $\epsilon$ yields the stated confidence bound. ∎
@@
$\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\zeta_{h}^{(W_{t}),t}(s,a),$
where $\zeta_{h}^{(W_{t}),t}(s,a)=\min\left\{H,\sqrt{\frac{4H^{2}\ln(mSAHT/\delta)}{\max\{1,N_{h}^{(W_{t}),t}(s,a)\}}}\right\}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{(W_{t}),t}(s,a)\right]\geq 1-\delta^{\prime}.$
@@
$\mathbb{P}\left[\bigcap_{i,h,t,s,a}\left\{\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\iota_{h}^{(W_{t}),t}(s,a)\right\}\right]\geq 1-mSAHT\delta^{\prime}.$
@@
[…]
@@ Appendix D Success Event
$F_{t}^{\alpha}$ :=$:=\left\{\exists s,a,h,i:\left|\hat{\alpha}_{i,h}^{t-1}(s,a)-\alpha_{i,h}(s,a)\right|\geq\zeta_{h}^{t-1}(s,a)\right\},$
$F_{t}^{r}$ :=$:=\left\{\exists s,a,h:\left|\hat{r}_{h}^{t-1}(s,a)-r_{h}(s,a)\right|\geq\phi_{h}^{r,t-1}(s,a)\right\},$
$F_{t}^{g}$ :=$:=\left\{\exists s,a,h,i:\left|\hat{g}_{i,h}^{t-1}(s,a)-g_{i,h}(s,a)\right|\geq\phi_{i,h}^{g,t-1}(s,a)\right\},$
$F_{t}^{p}$ :=$:=\left\{\exists s,a,s^{\prime},h:\left\|p_{h}(s^{\prime}\!\mid\!s,a)-\hat{p}_{h}^{t-1}(s^{\prime}\!\mid\!s,a)\right\|_{1}H\geq\phi_{h}^{p,t-1}(s,a,s^{\prime})\right\},$
$F_{t}^{N}$ :=$:=\left\{\exists s,a,h:N_{h}^{t-1}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right)\right\}.$
@@
  • Using concentration inequalities for multinomial distributions [21] and the union bound, we derive $\mathbb{P}[F^{p}]\leq\delta^{\prime}$.
  • Employing similar techniques as in [8], by bounding occupancy measure deviations, we obtain $\mathbb{P}[F^{N}]\leq\delta^{\prime}$.
@@ Appendix E Policy Evaluation
Algorithm 3 TPE (Truncated Policy Evaluation [11])
@@
5:         $\hat{Q}_{h}^{\pi_{t}}(s,a;\overline{g}_{i},\overline{p})=\min\Big{\{}\overline{g}_{i,h}(s,a)+\overline{p}_{h}(\cdot\mid s,a)\hat{V}_{\overline{g},h+1}^{\pi_{t}}(\cdot),H-h+1\Big{\}}$  ($\forall i\in[m]$)
@@ Appendix F Feasibility
To prove the lemma we show that under the success event $\mathcal{E}$, which holds with probability at least $1-\delta$, Projection (4) admits a feasible solution $q^{\pi_{1}}$. Precisely, under the success event $\mathcal{E}$, we have, for any feasible solution $q^{\pi_{1}}$ under Assumption 2 and any episode $t\in[T]$,
$\overline{G}^{\top}q^{\pi_{1}}\succeq G_{t}^{\top}q^{\pi_{1}}\succeq\alpha+\zeta_{t}=\overline{\alpha}\succeq\alpha,$
where the first inequality holds by the optimism of the constraints. The second and third inequalities show that if $q^{\pi_{1}}$ satisfies the constraints with respect to the true mean constraint matrix, it also satisfies the pessimistic thresholds. Thus, the feasible solutions are all available at every episode. ∎
@@
$\overline{G}^{\top}q^{\pi_{2}}\succeq G_{t}^{\top}q^{\pi_{2}}\succeq\alpha\succeq\alpha-\zeta_{t}=\underline{\alpha},$
@@ Appendix G Optimality and Optimism of Policies
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to time step $h$ in episode $t$. Suppose that for every $(t,h)$, the term $\zeta_{h}^{t}(s_{h},a_{h})$ satisfies
$\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})\!\mid\!\mathcal{F}_{h-1}^{t}]\right|\leq 2H.$
Then, for any confidence parameter $\delta\in(0,1)$, with probability at least $1-\delta$,
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\left(\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\!\mid\!\mathcal{F}_{h-1}^{t}\right]\right)\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right).$
@@
$k(t,h)=(t-1)H+h$
for $t\in[T]$ and $h\in[H]$ and relabel the filtration by $\mathcal{F}_{k}=\mathcal{F}_{h}^{t}$ and $\mathcal{F}_{0}:=\{\emptyset,\Omega\}$. For each $k\in[TH]$, set
$D_{k}:=\zeta_{k}(s_{h},a_{h})-\mathbb{E}[\zeta_{k}(s_{h},a_{h})\mid\mathcal{F}_{k-1}].$
which holds that $\mathcal{F}_{k^{\prime}}\subset\mathcal{F}_{k}$ for any $k^{\prime}\leq k$.
For any $k\in[TH]$,
$\mathbb{E}[D_{k}\mid\mathcal{F}_{k-1}]=0,$
which means that $\{D_{k}\}_{k}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{k-1}\}$ and satisfies $|D_{k}|\leq 2H$. By the Azuma-Hoeffding inequality, for any $\epsilon>0$,
$\mathbb{P}\left(\left|\sum_{k=1}^{T\cdot H}D_{k}\right|>\epsilon\right)\leq 2\text{exp}\left(-\frac{\epsilon^{2}}{2TH\cdot(2H)^{2}}\right)=2\text{exp}\left(-\frac{\epsilon^{2}}{8TH^{3}}\right).$
Setting the right-hand side equal to $\delta$ yields $\epsilon=2\sqrt{2TH^{3}\text{ln}\frac{2}{\delta}}$. Then, with probability at least $1-\delta$, we have
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\left(\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid\mathcal{F}_{h-1}^{t}\right]\right)\right|=\left|\sum_{k=1}^{T\cdot H}D_{k}\right|\leq 2\sqrt{2TH^{3}\text{ln}\frac{2}{\delta}}\lesssim\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}.$
@@
$N_{h}^{(W_{t}),t}(s,a)>\frac{1}{2}\sum_{j=\max\{1,t-W_{t}+1\}}^{t}q_{h}^{\pi_{t}}(s,a;p)-H\ln\frac{SAH}{\delta^{\prime}},$
@@
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t}(s,a),1\}}}\mid\mathcal{F}_{t}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T/\gamma}+SAH).$
@@
We first define $w_{h}^{t}(s,a)=\mathbb{P}\{s_{h}^{t}=s,a_{h}^{t}=a\}$. Then, we have
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{{N_{h}^{(W_{t}),t}(s,a),1\}}}}\mid\mathcal{F}_{t}\right]$ $=\sum_{t,h,s,a}w_{h}^{t}(s,a)\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t}(s,a),1\}}}$
$\leq\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}\sqrt{\frac{1}{N_{h}^{(W_{t}),t}(s,a)}}+\sum_{t,h}\sum_{s,a\in L_{t}^{c}}w_{h}^{t}$
$\leq\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}\sqrt{\frac{1}{N_{h}^{(W_{t}),t}(s,a)}}+SAH,$
where we set $L_{t}=\left\{N_{h}^{(W_{t}),t}(s,a)\geq 1\right\}$ and thus set $L_{t}^{c}=\left\{N_{h}^{(W_{t}),t}(s,a)=0\right\}$. Consider the following term,
$\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}\sqrt{\frac{1}{N_{h}^{(W_{t}),t}(s,a)}}$ $\leq\sqrt{\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}}\sqrt{\sum_{t,h}\sum_{s,a\in L_{t}}\frac{w_{h}^{t}}{N_{h}^{(W_{t}),t}(s,a)}}$
$\leq\sqrt{\sum_{t,h,s,a}w_{h}^{t}}\sqrt{\sum_{t,h}\sum_{s,a\in L_{t}}\frac{w_{h}^{t}}{N_{h}^{(W_{t}),t}(s,a)}}$
$\leq\sqrt{TH}\sqrt{\sum_{t,h,s,a}\frac{w_{h}^{t}}{N_{h}^{(W_{t}),t}(s,a)}}\leq\tilde{O}(\sqrt{SAH^{2}T/\gamma}).$

### v3

@@ Appendix C Estimators of Thresholds
For every constraint $i\in[m]$, state $s$, action $a$, step $h\in[H]$ and episode $t\in[T]$, let $(s_{h}^{l},a_{h}^{l})$ denotes the state-action pair visited in episode $l$ at step $h$; $(s_{h}^{l},a_{h}^{l},s_{h+1}^{l})$ represents the state-action pair $(s_{h}^{l},a_{h}^{l})$ is visited and the environment evolves to next state $s_{h+1}^{l}$ at step $h$ in episode $l$; $\mathbf{1}_{X}$ is the indicator function of $X$; and $N_{h}^{t}(s,a)=\sum_{l=1}^{t-1}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ is the total number of visits to the pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ at step $h$ before episode $t\in[T]$. First, we give the empirical averages of the thresholds as follows:
$\hat{\alpha}_{i,h}^{t}(s,a):=\frac{\sum_{l=1}^{t}\tilde{\alpha}_{i,h}^{l}(s,a)\,\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}}{\max\{1,N_{h}^{t}(s,a)\}}.$
@@
$\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{t}(s,a),$
where $\iota_{h}^{t}(s,a):=\sqrt{\frac{\ln\left(\frac{2}{\delta}\right)}{2N_{t}(s,a)}}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\geq\frac{c}{N_{t}(s,a)}\right]\leq 2\exp\left(-\frac{2c^{2}}{N_{t}(s,a)}\right).$
Setting $\delta=2\exp\left(-\frac{2c^{2}}{N_{t}(s,a)}\right)$ and solving to find a proper value of $c$ gives the result for the threshold function. ∎
@@
$\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\zeta_{h}^{t}(s,a),$
where $\zeta_{h}^{t}(s,a):=\min\left\{1,\sqrt{\frac{4\ln(mSAHT/\delta)}{\max\{1,N_{t}(s,a)\}}}\right\}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{t}(s,a)\right]\geq 1-\delta^{\prime}.$
@@
$\mathbb{P}\left[\bigcap_{s,a,m,h,t}\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{t}(s,a)\right\}\right].$
@@
$\mathbb{P}\left[\bigcap_{s,a,m,h,t}\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{t}(s,a)\right\}\right]$
$=$ $1-\mathbb{P}\left[\bigcup_{s,a,m,h,t}\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{t}(s,a)\right\}^{c}\right]$
$=$ $1-\sum_{s,a,m,h,t}\mathbb{P}\left[\left\{\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{t}(s,a)\right\}^{c}\right]$
@@
where the inequality holds by union bound. Noticing that $\alpha_{i,h}^{t}(s,a)\leq H$, we substitute $\delta^{\prime}$ with $\delta:=\frac{\delta^{\prime}}{mSAHT}$ in $\iota_{t}(s,a)$ with an additional union bound over the possible values of $N_{t}(s,a)$. This yields $\zeta_{t}(s,a)$, and concludes the proof. ∎
@@
$\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{(W_{t}),t}(s,a),$
where $\iota_{h}^{(W_{t}),t}(s,a)=\sqrt{\frac{\ln\left(\frac{2}{\delta}\right)}{2\max\{1,N_{h}^{(W_{t}),t-1}(s,a)\}}}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}\right|\geq\epsilon\right]\leq 2\exp\left(-2\epsilon^{2}N_{h}^{(W_{t}),t-1}(s,a)\right).$
Setting $\delta=2\exp\left(-2\epsilon^{2}N_{h}^{(W_{t}),t-1}(s,a)\right)$ and solving for $\epsilon$ yields the stated confidence bound. ∎
@@
$\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}\right|\leq\zeta_{h}^{(W_{t}),t}(s,a),$
where $\zeta_{h}^{(W_{t}),t}(s,a)=\min\left\{1,\sqrt{\frac{4\ln(mSAHT/\delta)}{\max\{1,N_{h}^{(W_{t}),t-1}(s,a)\}}}\right\}$.
@@
$\mathbb{P}\left[\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{(W_{t}),t}(s,a)\right]\geq 1-\delta^{\prime}.$
@@
$\mathbb{P}\left[\bigcap_{i,h,t,s,a}\left\{\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}\right|\leq\iota_{h}^{(W_{t}),t}(s,a)\right\}\right]\geq 1-mSAHT\delta^{\prime}.$
@@
$\hat{r}_{h}^{t}(s,a)$ $:=\frac{\sum_{l=1}^{t}\tilde{r}_{h}^{l}(s,a)\,\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}}{\max\{1,N_{h}^{t}(s,a)\}},$ (6a)
$\hat{g}_{i,h}^{t}(s,a)$ $:=\frac{\sum_{l=1}^{t}\tilde{g}_{i,h}^{l}(s,a)\,\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}}{\max\{1,N_{h}^{t}(s,a)\}},$ (6b)
@@ Appendix D Success Event
$F_{t}^{\alpha}$ $:=\left\{\exists s,a,h,i:\left|\hat{\alpha}_{i,h}^{t}(s,a)-\alpha_{i,h}\right|\geq\zeta_{h}^{t}(s,a)\right\},$
$F_{t}^{r}$ $:=\left\{\exists s,a,h:\left|\hat{r}_{h}^{t}(s,a)-r_{h}(s,a)\right|\geq\phi_{h}^{r,t}(s,a)\right\},$
$F_{t}^{g}$ $:=\left\{\exists s,a,h,i:\left|\hat{g}_{i,h}^{t}(s,a)-g_{i,h}(s,a)\right|\geq\phi_{i,h}^{g,t}(s,a)\right\},$
$F_{t}^{p}$ $:=\left\{\exists s,a,s^{\prime},h:\left|p_{h}(s^{\prime}\mid s,a)-\hat{p}_{h}^{t-1}(s^{\prime}\mid s,a)\right|\geq\phi_{h}^{p,t}(s,a,s^{\prime})\right\},$
$F_{t}^{N}$ $:=\left\{\exists s,a,h:N_{h}^{t-1}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right)\right\}.$
@@
  • Using concentration inequalities for multinomial distributions [maurer2009empirical] and the union bound, we derive $\mathbb{P}[F^{p}]\leq\delta^{\prime}$.
  • Employing similar techniques as in [dann2017unifying], by bounding occupancy measure deviations, we obtain $\mathbb{P}[F^{N}]\leq\delta^{\prime}$.
@@ Appendix E Policy Evaluation
Algorithm 3 TPE (Truncated Policy Evaluation [efroni2020exploration])
@@
5:    $\hat{Q}_{h}^{\pi_{t}}(s,a;\overline{g}_{i},\overline{p})=\min\Big\{\overline{g}_{i,h}(s,a)+\overline{p}_{h}(\cdot\mid s,a)\hat{V}_{\overline{g},h+1}^{\pi}(\cdot),H-h+1\Big\}$  ($\forall i\in[m]$)
@@ Appendix F Feasibility
To prove the lemma we show that under the success event $\mathcal{E}$, which holds with probability at least $1-\delta$, Projection (3) admits a feasible solution $q^{\pi^{0}}$. Precisely, under the success event $\mathcal{E}$, we have, for any feasible solution $q^{\pi^{0}}$ under Assumptions 1 and 2 and any episode $t\in[T]$,
$\overline{G}^{\top}q^{\pi^{0}}\succeq G_{t}^{\top}q^{\pi^{0}}\succeq\alpha+\zeta_{t}=\overline{\alpha}\succeq\alpha,$
where the first inequality holds by the optimism of the constraints. The second and third inequalities show that if $q^{\pi^{0}}$ satisfies the constraints with respect to the true mean constraint matrix, it also satisfies the pessimistic thresholds. Thus, the feasible solutions are all available at every episode. ∎
@@
$\overline{G}^{\top}q^{\pi^{0}}\succeq G_{t}^{\top}q^{\pi^{0}}\succeq\alpha\succeq\alpha-\zeta_{t}=\underline{\alpha},$
@@ Appendix G Optimality and Optimism of Policies
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to time step $h$ in episode $t$. Suppose that for every $(t,h)$, $\zeta_{h}^{t}(s_{h},a_{h})$ satisfies
$\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})]\mid\mathcal{F}_{h}^{t-1}\right|\leq 2.$
Then, with probability $1-\delta$,
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\left(\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid\mathcal{F}_{h}^{t-1}\right]\right)\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH\text{ln}\frac{1}{\delta}}\right).$
@@
k(t,h)=(t1)H+hk(t,h)=(t-1)H+h
for $t\in[T]$ and $h\in[H]$. For each $k\in[TH]$, set
$D_{k}:=\zeta_{k}(s_{h},a_{h})-\mathbb{E}[\zeta_{k}(s_{h},a_{h})\mid\mathcal{F}_{h}^{t-1}].$
which holds that $\mathcal{F}_{h}^{t}\subset\mathcal{F}_{h^{\prime}}^{t^{\prime}}$ for any $t\leq t^{\prime}$.
For any $k(t,h)\in[T]\times[H]$,
$\mathbb{E}[D_{k}\mid\mathcal{F}_{h}^{t-1}]=0,$
which means that $\{D_{k}\}_{k}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{h}^{t-1}\}$ and satisfies $|D_{k}|\leq 2$. By the Azuma-Hoeffding inequality, for any $\epsilon>0$,
$\mathbb{P}\left(\sum_{k=1}^{T\cdot H}D_{k}>\epsilon\right)\leq 2\text{exp}\left(-\frac{\epsilon^{2}}{2TH\cdot(2)^{2}}\right)=2\text{exp}\left(-\frac{\epsilon^{2}}{8TH}\right).$
Setting the right-hand side equal to $\delta$ yields $\epsilon=2\sqrt{2TH\text{ln}\frac{2}{\delta}}$. Then, with probability $1-\delta$, we have
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid\mathcal{F}_{h}^{t-1}\right]\right|=\left|\sum_{k=1}^{T\cdot H}D_{k}\right|\leq 2\sqrt{2TH\text{ln}\frac{2}{\delta}}\lesssim\sqrt{TH\text{ln}\frac{1}{\delta}}.$
@@
$N_{h}^{(W_{t}),t-1}(s,a)>\frac{1}{2}\sum_{j=\max\{1,t-W_{t}+1\}}^{t-1}q_{h}^{\pi_{t}}(s,a;p)-H\ln\frac{SAH}{\delta^{\prime}},$
@@
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t-1}(s,a),1\}}}\mid\mathcal{F}_{t-1}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T/\gamma}+SAH).$
@@
We first define $w_{h}^{t}(s,a)=\mathbb{P}(s_{h}^{t}=s,a_{h}^{t}=a|\mathcal{F}_{t-1})$. Then, we have
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{\max\{{N_{h}^{(W_{t}),t-1}(s,a),1\}}}}\mid\mathcal{F}_{t-1}\right]$ $=\sum_{t,h,s,a}w_{h}^{t}(s,a)\sqrt{\frac{1}{\max\{N_{h}^{(W_{t}),t-1}(s,a),1\}}}$
$\leq\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}\sqrt{\frac{1}{N_{h}^{(W_{t}),t-1}(s,a)}}+\sum_{t,h}\sum_{s,a\in L_{t}^{c}}w_{h}^{t}$
$\leq\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}\sqrt{\frac{1}{N_{h}^{(W_{t}),t-1}(s,a)}}+SAH,$
where we set $L_{t}=\left\{N_{h}^{(W_{t}),t-1}(s,a)\geq 1\right\}$ and thus set $L_{t}^{c}=\left\{N_{h}^{(W_{t}),t-1}(s,a)=0\right\}$. Consider the following term,
$\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}\sqrt{\frac{1}{N_{h}^{(W_{t}),t-1}(s,a)}}$ $\leq\sqrt{\sum_{t,h}\sum_{s,a\in L_{t}}w_{h}^{t}}\sqrt{\sum_{t,h}\sum_{s,a\in L_{t}}\frac{w_{h}^{t}}{N_{h}^{(W_{t}),t-1}(s,a)}}$
$\leq\sqrt{\sum_{t,h,s,a}w_{h}^{t}}\sqrt{\sum_{t,h}\sum_{s,a\in L_{t}}\frac{w_{h}^{t}}{N_{h}^{(W_{t}),t-1}(s,a)}}$
$\leq\sqrt{TH}\sqrt{\sum_{t,h,s,a}\frac{w_{h}^{t}}{N_{h}^{(W_{t}),t-1}(s,a)}}\leq\tilde{O}(\sqrt{SAH^{2}T/\gamma}).$